### v1

# Zero Indirect Band Gap in Non-Hermitian Systems

Abstract
Zero indirect gaps in band models are typically viewed as unstable and achievable only through fine-tuning. Recent works, however, have revealed robust semimetallic phases in Hermitian systems where the indirect gap remains pinned at zero over a finite parameter range. Here, we extend this paradigm to non-Hermitian lattice models by studying a one-dimensional diamond-like system with gain and loss. We show that a zero indirect band gap can remain stable against non-Hermitian perturbations and identify the regimes where this robustness persists. Remarkably, we find that the zero indirect gap induces a suppression of the non-Hermitian skin effect distinct from other physical mechanics already discussed in the literature. Our results reveal new connections between indirect gaps, exceptional points and non-Hermitian skin effect, opening avenues for experimental realizations.
I Introduction
The distinction between direct and indirect band gaps plays a central role in determining the electronic and optical properties of condensed matter systems [1, 2, 3]. While the direct band gap occurs at the same momentum points, the indirect band gap arises when the minimum of the conduction band and maximum of valence band occurs at different momenta. In this context, a zero indirect band gap is usually thought as unstable and induced by a mere fine tuning of the physical parameters of the systems. However, it has been recently shown the existence of suitable one- [4] and two-dimensional Hermitian lattice models [5, 6, 7, 8], in which the zero indirect band gap is robust and survives for a finite range of values of the physical parameters. This allows us to define of a novel family of semimetals in which bands do not touch each other differently from the more conventional semimetallic phases such as topological [9, 10, 11, 12, 13] and geometric semimetals [14] that instead support Dirac-like and Weyl-like cones. Importantly, one of these semimetallic phases with zero indirect band gap and non-trivial Chern number, dubbed two-dimensional Chern semimetal [5], has been experimentally realized in photonic crystals [15]. Thus, it is then relevant at both theoretical and experimental level to figure out the behavior of these semimetallic phases in the presence of dissipation (gain and loss) and possibly show the existence of robust indirect band gaps in non-Hermitian phases.
We remind that in recent years, condensed and synthetic matter systems have been revisited from the perspective of non-Hermitian systems [16], in which non-Hermitian Hamiltonians in open quantum systems emerge due to the presence of gain and loss. Here, the corresponding energy spectra can acquire complex values due to the nature of non-Hermiticity of the Hamiltonians. Non-Hermitian systems can exhibit distinctive features such as exceptional points (EPs) [17, 18, 19, 20, 21], the non-Hermitian skin effect (NHSE) [22, 23, 24, 25, 26], and non-Hermitian topology [27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. In this framework, it is particularly intriguing to explore how zero indirect band gaps manifest within the complex band structures of non-Hermitian systems and how they influence the presence of EPs and the NHSE. Despite their significance, systematic studies at the intersection of non-Hermitian physics and indirect band gaps have been completely overlooked, providing a key motivation for the present work.
The main goal of this paper is to show the existence of robust zero indirect band gaps in non-Hermitian models. For simplicity, we will focus on an one-dimensional system on a diamond-like lattice and study the presence of EPs and the NHSE. By systematically varying the parameters of our non-Hermitian model, we will identify the ranges where the indirect band gap remains pinned at zero, thereby marking its stability. Importantly, we will show a suppression of the NHSE due to the zero indirect gap, which represents one of the main results of our work. Notice that this NHSE suppression is genuinely distinct from other mechanisms induced by magnetic and electric fields and by anyonic statistics [37, 38, 39, 40]. Thus, our work paves the way for the exploration of new directions in non-Hermitian band theory with relevant implications for the experimental realization of semimetallic phases with robust zero indirect band gaps.
II Model Hamiltonian
We start considering a one dimensional system on a diamond-like chain with three sites per unit cell, which generalizes the Hermitian system introduced in Ref. [4]. The tight-binding Hamiltonian is written in terms of fermionic basis as
$H$ $=\sum_{i}J\big(a^{\dagger}_{i+1}b_{i}+b^{\dagger}_{i+1}c_{i}+\text{h.c.}\big)-h\big(a^{\dagger}_{i}b_{i}-b^{\dagger}_{i}c_{i}+\text{h.c.}\big)$
$\quad+\beta_{1}a^{\dagger}_{i}c_{i}+\beta_{2}c^{\dagger}_{i}a_{i}$ (1)
where J and h are real parameters. $\beta_{1}$ and $\beta_{2}$ are generic complex numbers, which take into account gain and loss between $a$ and $c$ sites in the same unit cell (on-site dissipation). The corresponding momentum space Hamiltonian is given by,
$H(k)=\begin{pmatrix}0&-h+Je^{-ik}&\beta_{1}\\ -h+Je^{ik}&0&-h+Je^{-ik}\\ \beta_{2}&-h+Je^{ik}&0\end{pmatrix}$.
Through out the manuscript J and h are set to 1 for simplicity. Despite this choice, it is possible to show that the zero indirect gap persists even for other values of the hopping parameters. Notice that this Hamiltonian does not have analytical spectrum except for $\beta_{1}=\beta_{2}$ = 0, for which the system is Hermitian and the spectrum exhibits a perfect flat middle band and two linear dispersive bands, such that all three bands touch each other at a Dirac-like point at $k=0$.
Figure 1: (a)-(d) Real component of the dispersion spectrum plotted as a function of k. Panels (a)-(d) correspond to distinct set of imaginary part of $Im(\beta_{1})=Im(\beta_{2})=0.2,0.8,1.52,2.0$ respectively. (e) Zero indirect gap plotted with respect to the $Im(\beta_{1})=Im(\beta_{2})$ for fixed value of $Re(\beta_{1})=Re(\beta_{2})$=1.
The momentum–space dispersion, shown in Fig. 1(a)-(c), reveals the emergence of a zero indirect band gap. This feature persists across different values of the imaginary components of $\beta_{1}$ and $\beta_{2}$, while keeping their real parts fixed at $Re(\beta_{1})=Re(\beta_{2})=1$. An horizontal red dotted line serves as a marker for zero indirect band gap. From Fig. 1(d), it is evident that as the imaginary parts of $\beta_{1}$ and $\beta_{2}$ increase, the system undergoes a qualitative change in its band structure. More specifically the zero indirect band gap gradually disappears, giving rise to a direct gap closure.
Figure 2: Parameter space plotted with respect to $Re(\beta_{1})$ and $Re(\beta_{2})$ for different values of imaginary components i.e., (a)-(d) $Im(\beta_{1})=Im(\beta_{2})=0.2,0.6,1.2,1.8$ respectively.
This signifies that the enhancement in the strength of non-Hermitian coupling, destabilizes the zero indirect band gap. The range of $Im(\beta_{1})=Im(\beta_{2})\equiv Im(\beta)$ for which the zero indirect band gap exist is presented in Fig.1(e) and is given by $1.52<Im(\beta)<1.52$. The existence of this finite range demonstrates that the zero indirect bang gap is indeed robust. To further investigate this regime, we map out the parameter space in terms of the real components of $\beta_{1}$ and $\beta_{2}$. In the Fig.2, we illustrate how a zero indirect band gap emerges within certain regions of the parameter space. This regime is analyzed for various values of $Im(\beta)$ indicating that it exists under the condition $Re(\beta_{1})=Re(\beta_{2})$.
III Exceptional Points
In this section, we investigate the appearance of EPs due to non-Hermitian coupling terms in our model. EPs arise in the parameter space where all the three bands touch each other. Thus, in this specific regime, we compute the condition number $\rm cond(V)$} [41, 42, 43]. In non-Hermitian systems, this number identifies the non-orthogonality of eigenvectors, which can become nearly linearly dependent.
The right eigenvectors of the Hamiltonian $H(\mathbf{k})$ are obtained by solving
$H(\mathbf{k})\,|R_{n}(\mathbf{k})\rangle=E_{n}(\mathbf{k})\,|R_{n}(\mathbf{k})\rangle,$ (2)
where $E_{n}(\mathbf{k})$ are (generally complex) eigenvalues. Writing these right eigenvectors as columns of a matrix $V(k)$,
$V(\mathbf{k})=\big(|R_{1}(\mathbf{k})\rangle,|R_{2}(\mathbf{k})\rangle,\dots,|R_{N}(\mathbf{k})\rangle\big),$ (3)
the condition number of $V(k)$ is defined as
$\mathrm{cond}\big(V)=\|V(\mathbf{k})\|\,\|V(\mathbf{k})^{-1}\|.$ (4)
Figure 3: Real component of the dispersion spectrum plotted with respect to k. The red dotted line corresponds to the profile of $\rm cond(V)$} plotted over k for the values $\beta_{1}=-0.5+0.5i$ and $\beta_{2}=0.5+0.5i$.
At EPs, not only two or more eigenvalues coalesce, but their associated eigenvectors also become linearly dependent. In this case, the matrix $V(\mathbf{k})$ becomes nearly singular, and hence ideally the condition number diverges to infinite. In Fig. 3, we present the behavior of the condition number $\rm cond(V)$} superimposed on the energy dispersion spectrum. The figure clearly illustrates that the sharp divergence of $\rm cond(V)$} occurs precisely at the points in momentum space where the eigenvalue bands coalesce. At these points, not only the eigenvalue bands but also the corresponding eigenvectors coalesce. This feature is indicated by the divergence of $\rm cond(V)$}.
Figure 4: Scaling behavior of the maximum condition number of the matrix V(k) as a function of the number of momentum points $N_{k}$ for different parameter choices of $\beta_{1}$ and $\beta_{2}$, namely $\beta_{1}=\{-0.5,1.0\},\;\beta_{2}=\{0.5,1.0\}$; $\beta_{1}=\{0.25,1.0\},\;\beta_{2}=\{-0.25,1.0\}$; and $\beta_{1}=\{0.5,1.0\},\;\beta_{2}=\{-0.5,1.0\}$ respectively.
This divergence serves as a marker of EPs, highlighting the loss of linear independence of the eigenvectors and the consequent ill-conditioning of the eigenvector matrix $V(k)$. Moreover to understand if the maximum value of $\rm cond(V)$} scales with the resolution of k, we do a systematic scaling analysis which is presented in Fig.4. Across all three cases, the scaling law follows an approximate power-law behavior of the form,
$\text{maximum cond}(V)\propto N_{k}^{\alpha},$ (5)
with a fitted exponent $\alpha=0.48$. The consistency of this exponent across different parameter regimes highlights the robustness of the scaling behavior. This robustness not only provides a reliable quantitative signature for the presence of EPs but also suggests a universal mechanism governing the sensitivity of eigenvectors near non-Hermitian degeneracies in our model.
IV Non-Hermitian Skin effect
Before reviewing the NHSE in our model, it is useful to briefly discuss the central role of NHSE in non-Hermitian systems. This peculiar effect refers to the significant number of bulk eigenstates accumulate at the boundary under open boundary conditions [24, 25, 26]. Physically the NHSE arises due to the asymmetric couplings or gain/loss [44] in the system. The existence of non-Hermiticity in both quantum materials and synthetic matter systems makes the NHSE very relevant from the perspective of experimental realizations [45, 46, 47]. Here, to characterize this effect induced by non-Hermitian couplings $\beta_{1}$ and $\beta_{2}$ in our model, we compute the average inverse participation ratio (IPR) of the eigenstates across the parameter space spanned by Re($\beta_{1}$) and Re($\beta_{2}$). The results reveal a clear suppression of NHSE along the diagonal line Re($\beta_{1}$) = Re($\beta_{2}$), where the non-Hermitian coupling is effectively symmetric and hence the asymmetry vanishes. In this regime the eigenstates are more delocalized, leading to smaller IPR values. In contrast, away from the diagonal, the asymmetry between Re($\beta_{1}$) and Re($\beta_{2}$) becomes significant, and the eigenstates exhibit strong localization reflected in the higher IPR values.
Figure 5: Average IPR of eigenstates as a function of $\mathrm{Re}(\beta_{1})$ and $\mathrm{Re}(\beta_{2})$ for different values of the imaginary component $\mathrm{Im}(\beta)$ (a) $0.2$, (b) $0.6$, (c) $1.2$, and (d) $1.8$.
By comparing Fig.2 and Fig.5, we notice a crucial connection between NHSE and existence of zero indirect band gap in the same parameter space. As shown previously, the zero indirect band gap dominates over the diagonal region. Remarkably, in this same regime, the IPR of the eigenstates remains small, which is a clear signature of delocalization. In other words, the NHSE, where the bulk eigenstates are accumulated of at the system’s edges, is significantly weakened due to the presence of the zero indirect gap. It would be relevant to check the weakening of the NHSE also in higher-dimensional gapless systems with zero indirect gaps. We leave this important point to future work.
V Conclusion
In this work, we have investigated the emergence of a robust zero indirect band gap in a non-Hermitian diamond-like chain with generic complex intra-cell couplings. By analyzing the momentum space dispersion spectrum, we have identified the parameter regime where the zero indirect band gap persists. The non-Hermitian character of the model Hamiltonian leads to rich phenomena such as the emergence of EPs, NHSE. We have computed condition number of the eigenvector matrix to identify the EPs and have obtained a scaling law that proves the divergence of condition number at the EPs. Strikingly, our results reveal that the zero indirect band gap can suppress the NSHE, thereby allowing the extended states to prevail. This interplay between zero indirect band gap and the non-Hermitian features provides new insights into spectral properties of non-Hermitian systems and supports in engineer novel lattice systems. As a future direction, we plan to investigate non-Hermitian phases in topological semimetals exhibiting robust zero indirect gaps in two [5] and higher dimensions, aiming to uncover novel spectral and transport phenomena arising from the interplay between topology and dissipation.